OpenStudy (anonymous):

The electrical resistance of certain material is related to its temperature by; R=50(1+0.0053T+0.00001T^2) Ohms Where T is in degrees Celsius.If the temperature is initially at 80°C and changes to 85°C,by how much,approximately,will the resistance change ??

OpenStudy (anonymous):

Ok First you want to find in general what delta R is. You are given a formula that you plug T.initial and T.final into. So R.initial = 50(1+0.0053(80) + 0.00001(80)^2) = 74.4 Ohms And R.final = 50(1+0.0053(85) + 0.00001(85)^2) = 76.1375 Ohms Delta R = R.final - R.initial = 76.1375 Ohms - 74.4 Ohms = 1.7375 Ohms Also, check the problem to see if your value should be positive or negative to make sure you plugged in R.initial and R.final in correctly or to avoid a math error. Since the temperature increased, and the resistance increases with temperature, a positive delta R value makes sense logically.